To double my money, if $600 is invested at 8% compounded monthly by choosing the appropriate formula and solving it for t? What is the formula?
Answer by stanbon(75887)

To double my money, if $600 is invested at 8% compounded monthly by choosing the appropriate formula and solving it for t? What is the formula?
-------
Standard compounding formula:: A(t) = P(1+(r/n))^(n*t)
---------
Doubling your investment gives:: 2*P = P(1+(r/n)^(n*t)
-------
Your Problem::
2*600 = 600(1+(0.08/12))^(12*t)
-------
(1.0067)^(12t) = 2
--------
12t*log(1.0067) = log(2)
------
12t = log(2)/log(1.0067) = 103.8
----
time = 103.8/12 = 8.65 years = 8yrs 8 months
-------------
Cheers,
Stan H.
-------------------